### v1

# Geometric Model Selection for Latent Space Network Models: Hypothesis Testing via Multidimensional Scaling and Resampling Techniques

Abstract
Latent space models assume that network ties are more likely between nodes that are closer together in an underlying latent space. Euclidean space is a popular choice for the underlying geometry, but hyperbolic geometry can mimic more realistic patterns of ties in complex networks. To identify the underlying geometry, past research has applied non-Euclidean extensions of multidimensional scaling (MDS) to the observed geodesic distances: the shortest path lengths between nodes. The difference in stress, a standard goodness-of-fit metric for MDS, across the geometries is then used to select a latent geometry with superior model fit (lower stress). The effectiveness of this method is assessed through simulations of latent space networks in Euclidean and hyperbolic geometries. To better account for uncertainty, we extend permutation-based hypothesis tests for MDS to the latent network setting. However, these tests do not incorporate any network structure. We propose a parametric bootstrap distribution of networks, conditioned on observed geodesic distances and the Gaussian Latent Position Model (GLPM). Our method extends the Davidson-MacKinnon J-test to latent space network models with differing latent geometries. We pay particular attention to large and sparse networks, and both the permutation test and the bootstrapping methods show an improvement in detecting the underlying geometry.
Keywords: network data, permutation test, bootstrapping, geodesic distance, hyperbolic geometry
1 Introduction
Networks are often modeled using latent space models where each node is assigned a position in a low-dimensional latent space (Hoff et al. [2002]; Krioukov et al. [2010]; Sweet and Adhikari [2020]; Sosa and Betancourt [2021]). The probability of having an edge between two nodes is proportional to the distance in the underlying space, so that pairs of nodes that are further apart are less likely to be connected. Euclidean space has been a popular default choice for the underlying space since the model was introduced by Hoff et al. [2002]. Yet there is a growing interest in alternative geometries, such as hyperbolic space.
Krioukov et al. [2010] argues that hyperbolic geometry is effective for complex networks: First, hyperbolic spaces expand faster than Euclidean spaces. Hyperbolic spaces expand exponentially, while Euclidean spaces expand polynomially. Many real-world networks tend to be treelike, and the structure of trees is similar to hyperbolic geometry in the sense that the number of nodes grows exponentially with the depth. Second, common network features, such as a skewed degree distribution and high amounts of clustering, are naturally generated from the basic properties of hyperbolic geometry. Furthermore, Smith et al. [2019] shows that modeling networks in hyperbolic space is promising, since it results in networks with higher levels of degree centrality, betweenness, and closeness without dramatic losses in clustering or average path length.
Current approaches to determining which geometry underlies a particular observed network include: (i) comparing the fit of multi-dimensional scaling (MDS) of geodesic network distances across different geometries, as in Papamichalis et al. [2022]; (ii) borrowing tools from metric spaces, such as $\delta$-hyperbolicity (Kennedy et al. [2013];Narayan and Saniee [2011]) or Van Der Hoorn et al. [2021]; and (iii) curvature estimation, via a clique-based latent distance estimator (Lubold et al. [2022]; Wilkins-Reeves and McCormick [2024]).
An MDS-based approach is appealing since it is metric-agnostic and requires minimal computational effort. We assess the effectiveness of the method proposed in Papamichalis et al. [2022] forgeometry selection through simulations of latent space networks in hyperbolic and Euclidean geometry. To make such geometry selection statistically principled, we further expand the approach to appropriately account for uncertainty. The standard way of accounting for statistical uncertainty in MDS is through permutation tests (Mair et al. [2016] as well as Mair et al. [2022]) or bootstrap-based methods (Mair et al. [2022]; Jacoby and Armstrong [2014]). However, these approaches are likely overly conservative, as they do not account for the type of structures we expect in network data. Our proposed method forms bootstrap distributions of network data, conditional on the observed shortest paths. Our approach builds on the success of using bootstrap distributions of networks to assess network structure Levin and Levina [2021].
In Section 2, we review MDS in hyperbolic and Euclidean spaces and show how comparisons of standard goodness-of-fit measures, stress and strain, can guide geometry choice. In Section 3, we introduce methods for quantifying uncertainty for MDS goodness-of-fit metrics, reviewing the permutation test and extending the bootstrap $J$-test to the network setting. In Section 4, we describe the latent space models in both hyperbolic and Euclidean space and demonstrate the efficiency and accuracy of the methods through simulations. Section 5 demonstrates the methods on several real-world datasets.
2 Multi-dimensional scaling in the network setting
Without loss of generality, we assume a network of size $N$ consists of a set of $N$ nodes $\mathcal{N}=\{n_{1},n_{2},\dots,n_{N}\}$. We refer to this network by an adjacency matrix $\mathbf{Y}=(Y_{ij})_{1\leq i,j\leq N}$. We focus on undirected, unweighted networks without self-loops, that is for $i,j=1,2,\dots N$
[…]
$Y_{ii}$ $=0$ (2)
Moreover, only connected networks are considered in this study; that is, all the elements in the geodesic distance matrix $\mathbf{D}(\mathbf{Y})$ are non-negative finite numbers.
Classical MDS ([Borg and Groenen, 2007, chap 12]), also known as principal coordinate analysis, uses observed dissimilarities to embed data in a lower-dimensional metric space. Given a dissimilarity matrix $\boldsymbol{\Delta}$, the goal is to find a coordinate matrix $\mathbf{X}_{\mathbb{M}}$ in a low-dimensional manifold $\mathbb{M}$ that minimizes a target function which measures stress or strain. Stress serves as a goodness-of-fit measure for MDS. Although multiple definitions of stress have been proposed, we use the standard definition given by,
$S_{\mathbb{M}}(\mathbf{Y})=\sqrt{\sum_{i,j}(\delta_{ij}-\hat{\delta}_{ij})^{2}}$
where $\mathbf{Y}$ is the adjacency matrix of the observed network, $\delta_{ij}$ is the measure of dissimilarity (e.g. geodesic distance, or shortest path lengths) between nodes $i$ and $j$, and $\hat{\delta}_{ij}$ is the distance between points $n_{i}$ and $n_{j}$ calculated in manifold $\mathbb{M}$ based on the coordinate matrix $\mathbf{X}_{\mathbb{M}}$ recovered by MDS.
2.1 Method 1: Observed Stress Difference
In Papamichalis et al. [2022], they compare the quality of multidimensional scaling (MDS) across different underlying geometries. The stress from classical and hyperbolic MDS are compared directly. The geodesic distance matrix is used as the dissimilarity matrix and if $S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y})<0$, then the hyperbolic space is preferred as the underlying latent space.
2.2 Method 2: Permutation Test
In the literature, both permutation tests and bootstrapping have been used to measure uncertainty associated with MDS goodness of fit metrics Mair et al. [2016], Mair et al. [2022], Farine and Carter [2022]
Following Mair et al. [2016], a permutation test for goodness of fit for MDS specifies null and alternative hypotheses as
$\mathbf{H_{0}}:$ Stress/configuration are obtained from a random permutation of dissimilarities.
$\mathbf{H_{A}}:$ Stress/configuration are obtained from something other than a random permutation of dissimilarities.
This null hypothesis is weakly informative and states that dissimilarities are exchangeable and that there is no structure beyond random chance. Translating these hypotheses to the comparison between MDS approaches and the latent space network setting, the null and alternative hypothesis are given as
$\mathbf{H_{0}}:$
Difference between hyperbolic and euclidean stresses/configurations is obtained as the difference between stress/configurations from a random permutation.
$\mathbf{H_{A}}:$
Difference between hyperbolic and euclidean stresses/configurations is obtained from something other than a random permutation.
Mair et al. [2016] provide two scenarios for setting up a permutation scheme:
S1: In the case of directly observed dissimilarities, the elements of $\boldsymbol{\Delta}$ can be permuted.
S2: For derived dissimilarities, they proposed a strategy for systematic column-wise permutations of the raw data.
Figure 1: Permutation schemes for classical MDS (left) and in a latent space network setting (right). On the right, the horizontal arrows represent the network generative process, where $\boldsymbol{Z}$ is a matrix of true underlying latent positions.
[…]
This method is described in the following:
Algorithm 1 Method 2: Permutation Test
1:for $i=1,2,\ldots,P$ do
2:  Permute $\mathbf{Y}$ and get the new permuted adjacency matrix $\tilde{\mathbf{Y}}_{i}$
3:  if New network with adjacency matrix $\tilde{\mathbf{Y}}_{i}$ is connected then
4:   Apply classical MDS and hydra on $\mathbf{D}(\tilde{\mathbf{Y}}_{i})$
5:   Compute and store stress difference $S_{\mathbb{H}^{2}}(\tilde{\mathbf{Y}}_{i})-S_{\mathbb{R}^{2}}(\tilde{\mathbf{Y}}_{i})$
6:  else
7:   Discard the network and go to next iteration
8:  end if
9:end for
10:Calculate $P(S_{\mathbb{H}^{2}}(\tilde{\mathbf{Y}})-S_{\mathbb{R}^{2}}(\tilde{\mathbf{Y}})\leq S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y}))$.
11:if $P(S_{\mathbb{H}^{2}}(\tilde{\mathbf{Y}})-S_{\mathbb{R}^{2}}(\tilde{\mathbf{Y}})\leq S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y}))<0.05$ then
12:  Reject the null and hyperbolic geometry is preferred
13:else
14:  Fail to reject the null and Euclidean geometry is preferred
15:end if
2.3 Method 3: Bootstrapping
Bootstrapping Networks
Levin and Levina [2021]
provide two methods to bootstrap a network. The first method is designed to generate bootstrap replicates of U-statistics, which does not apply to our case. For the second method, they generate bootstrap replicates of whole networks and then evaluate network quantities. They work with the random dot product graph(RDPG) in which the latent positions $X_{i}$ are i.i.d. with inner product distribution $F$. In order to obtain bootstrapping replicates, they apply adjacency spectral embedding(ASE) on the adjacency matrix $A$, and get estimated latent positions $\hat{X}_{i}$. Then, let $\hat{F}_{n}$ denote the empirical distribution of $\hat{X}_{i}$ and bootstrap sample $X_{i}^{*}$ are drawn i.i.d. from $\hat{F}_{n}$. Bootstrap samples generated in this way have adjacency matrices similar to the observed one.
In our setting, we apply MDS to the geodesic distance matrix and get a coordinate matrix, but this matrix is not the latent positions. To generate bootstrap samples with a geodesic distance matrix that have a distribution similar to the observed network, we try to associate the geodesic distance with the distance between two nodes. In this new bootstrapping method, we estimate the distribution of the distance between two nodes conditioned on their observed geodesic distance and generate the distance matrix based on it. Then, generate a new bootstrap adjacency matrix.
Distribution of geodesic distance between two nodes given their positions
Fronczak et al. [2004]
discuss an analytic solution for the average path length in a random graph, within which they provide a method to find the probability that the geodesic distance of $n_{i}$ and $n_{j}$ is exactly $k$. Rastelli et al. [2016] uses the same reasoning and finds explicit expression under Gaussian LPM. To summarize this method:
First, define $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as the probability of at least one walk of length $k$ between $n_{i}$ and $n_{j}$. A walk is called a path only when all the nodes it passes through are distinct, while a walk does not require all its nodes to be distinct. In other words, $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ could also be thought of as the probability that the geodesic distance between these two nodes is not larger than $k$.
Then, define $\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as the probability that the geodesic distance between $n_{i}$ and $n_{j}$ is exactly $k$. Given the definition of $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$, we have:
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})-p_{k-1}(\mathbf{z}_{i},\mathbf{z}_{j})$ (3)
We further define $\xi_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ to be the probability that there exits a walk of length $k$ from $n_{i}$ to $n_{j}$ that passes through nodes $\{n_{i},n_{A_{m}^{1}},\dots,n_{A_{m}^{k-1}},n_{j}\}$, and we could refer to this walk of $k+1$ nodes event $A_{m}$. By the definition of a walk, we do not require $n_{A_{m}^{1}},\dots,n_{A_{m}^{k-1}}$ to be distinct. The number of such events is about $N^{k-1}$. Let $A=\cup_{m=1}^{N^{k-1}}A_{m}$, where $A$ is the event that at least one walk of length $k$ between $n_{i}$ and $n_{j}$. By Lemma 1 of Fronczak et al. [2004], we have:
$p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})\approx 1-\exp\{-N^{k-1}\xi_{k}(\mathbf{z}_{i},\mathbf{z}_{j})\}$ (4)
Note that the assumption for that lemma is that all events $A_{1},A_{2},\dots$ need to be mutually independent. However, a same shorter walk could appear in multiple events; then, there is a correlation between these events. As Fronczak et al. [2004] points out, the proportion of such correlations is negligible for short walks($k\ll N$).
Moreover, Rastelli et al. [2016] has shown that the quantities above could be estimated using the following explicit expression:
$\begin{cases}h_{r+1}=h_{r}\alpha_{r}^{-d}(2\pi\phi)^{\frac{d}{2}}f_{d}\left(\mathbf{z}_{i};\mathbf{0},\frac{\omega_{r}+\gamma}{\alpha^{2}_{r}}\right)\\ \alpha_{r+1}=\frac{\alpha_{r}\gamma}{\omega_{r}+\gamma}\\ \omega_{r+1}=\frac{\omega_{r}\phi+\omega_{r}\gamma+\gamma\phi}{\omega_{r}+\gamma}\end{cases}\text{, where}\begin{cases}h_{1}=\tau(2\pi\phi)^{\frac{d}{2}}\\ \alpha_{1}=1\\ \omega_{1}=\phi\end{cases}$ (5)
$\xi_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ $=h_{k}f_{d}\left(\mathbf{z}_{j}-\alpha_{k}\mathbf{z}_{i};\mathbf{0},\omega_{k}\right)\text{, for }k=1,2,\dots,N-1$ (6)
$\begin{split}\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})&=p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})-p_{k-1}(\mathbf{z}_{i},\mathbf{z}_{j})\\ &=\exp\{-N^{k-2}\xi_{k-1}(\mathbf{z}_{i},\mathbf{z}_{j})\}-\exp\{-N^{k-1}\xi_{k}(\mathbf{z}_{i},\mathbf{z}_{j})\}\end{split}$ (7)
As suggested in Rastelli et al. [2016], when a network is given, setting $\gamma=1$, $\phi$ and $\tau$ could be found through ad-hoc method that matches the observed and theoretical value of average degree and clustering coefficient through equation (15) and (16).
Distribution of distance between nodes given their geodesic distance
By definition, we could write $\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as conditional probability, that is,
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|\mathbf{z}_{i},\mathbf{z}_{j})$ (8)
Based on the Gaussian LPM (14), we know that the probability of having an edge depend only on $d_{ij}=d_{\mathbb{R}^{d}}(\mathbf{z}_{i},\mathbf{z}_{j})$. Thus, we could also rewrite equation (8)
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|d_{ij})$ (9)
Then, the distribution of distance between nodes given their geodesic distance could be found through:
$P(d_{ij}|\delta_{ij}=k)\propto P(\delta_{ij}=k|d_{ij})P(d_{ij})$ (10)
When $d=2$ and $\gamma$ set to be $1$,
$d_{ij}$ $=|\mathbf{z}_{i}-\mathbf{z}_{j}|=\sqrt{(z_{i1}-z_{j1})^{2}+(z_{i2}-z_{j2})^{2}}$ (11)
$\frac{d_{ij}}{\sqrt{2}}$ $=\sqrt{\left(\frac{z_{i1}-z_{j1}}{\sqrt{2}}\right)^{2}+\left(\frac{z_{i2}-z_{j2}}{\sqrt{2}}\right)^{2}}$ (12)
Based on the Gaussian LPM (14), $z_{i1},z_{j1},z_{i2}$, and $z_{j2}$ are independent and identically distributed, and follow $N(0,1)$. Then $\frac{z_{i1}-z_{j1}}{\sqrt{2}}$ and $\frac{z_{i2}-z_{j2}}{\sqrt{2}}$ are independent and identically distributed and follow $N(0,1)$. By the definition of the Chi distribution, we know that $d_{ij}/\sqrt{2}$ follows the Chi distribution with degree of freedom $2$, and $P(d_{ij})$ could easily be obtained.
Davidson-MacKinnon $J$-test
$J$
-test is a classical way to compare two nonnested regression models, and the core idea is that if model $M_{2}$ captures the missing structure of model $M_{1}$, then $M_{2}$’s fitted value should add explanatory power. In Section 5 of Davidson and MacKinnon [2002], they discuss the use of parametric bootstrap in $J$ test, with hypotheses:
$H_{0}$: The competing model does not provide additional explanatory power beyond the current model.
$H_{1}$: The competing model provides additional explanatory power.
To start with, they find $J$ statistic for the observed data set, denoted by $\hat{J}$. Then, to perform the bootstrapping, they fit the data into one of the two models and find the parameter estimates. These estimated parameters could later be used in the bootstrap data generating process. Once they have $B$ bootstrap samples, $J$ statistics are calculated again for these samples using the same method as the observed data set and are denoted by $J_{j}^{*},j=1,\dots,B$. Finally, the p-value for this test is given by
[…]
In the case of the latent space model, we adopt this bootstrapping method of the Davidson-MacKinnon $J$ test. Similar to the Davidson-MacKinnon $J$-test, we are comparing two models: GLPM and hyperbolic geometry model. The hypotheses are:
$H_{0}$: The hyperbolic model does not provide additional explanatory power beyond the euclidean model.
$H_{1}$: The hyperbolic model provides additional explanatory power.
We first found the stress difference between classical MDS and hydra. Then, just as they estimate parameter to generate bootstrap data, by using the distribution of distance between nodes given geodesic distance, we could generate bootstrap networks according to GLPM. Then, calculate stress differences for these bootstrap networks and find the p-value. This method is described in the following:
Algorithm 2 Method 3: Bootstrapping
1:for $k=1,2,\dots K$ do
2:  Compute and store $P(d_{ij}|\delta_{ij}=k)$
3:end for
4:for $i=1,2,\ldots,B$ do
5:  Generate distance matrix $\tilde{\mathbf{d}}_{\mathbb{R}^{2}}=(\tilde{d}_{ij})_{1\leq i,j\leq N}$. $\tilde{d}_{ij}$ is drawn from $P(d_{ij}|\delta_{ij})$.
6:  Generate $\tilde{\mathbf{Y}}_{i}$ based on $\tilde{\mathbf{d}}_{\mathbb{R}^{2}}$ and GLPM
7:  if New network with adjacency matrix $\tilde{\mathbf{Y}}_{i}$ is connected then
8:   Apply classical MDS and hydra on $\mathbf{D}(\tilde{\mathbf{Y}}_{i})$
9:   Compute and store stress difference $S_{\mathbb{H}^{2}}(\tilde{\mathbf{Y}}_{i})-S_{\mathbb{R}^{2}}(\tilde{\mathbf{Y}}_{i})$
10:  else
11:   Discard the network and go to next iteration
12:  end if
13:end for
14:Calculate $P(S_{\mathbb{H}^{2}}(\tilde{\mathbf{Y}})-S_{\mathbb{R}^{2}}(\tilde{\mathbf{Y}})\leq S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y}))$.
15:if $P(S_{\mathbb{H}^{2}}(\tilde{\mathbf{Y}})-S_{\mathbb{R}^{2}}(\tilde{\mathbf{Y}})\leq S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y}))<0.05$ then
16:  Reject the null and hyperbolic geometry is preferred
17:else
18:  Fail to reject the null and Euclidean geometry is preferred
19:end if
3 Simulations
We want to show the efficiency of the methods mentioned in the previous section. To achieve this goal, we simulate networks in Euclidean and hyperbolic latent spaces, using the models described in the following sections.
3.1 Latent space models
Modeling assumptions for latent space model are as follows:
  • Each node $n_{i}$ has an unknown position $\mathbf{z}_{i}$ in $t$-dimensional latent space $\mathbb{L}^{t}$
  • Latent positions $\mathbf{z}_{i}$’s are independent and follow some distributions in $\mathbb{L}^{t}$
  • Edges are conditionally independent given latent positions and the probability of an edge between node $n_{i}$ and $n_{j}$ depends on a function $s_{\mathbb{L}^{t}}(\mathbf{z}_{i},\mathbf{z}_{j})$.
3.1.1 Euclidean Geometry
Most commonly mentioned latent space models for networks that use Euclidean geometry as underlying latent space is the ”Distance Models” proposed in Hoff et al. [2002]. Later, Rastelli et al. [2016] proposed the Gaussian latent position model(Gaussian LPM, GLPM), in which a non-normalized Gaussian density is used for edge instead of logistic link function in Distance Models. We will use the Gaussian LPM as the Euclidean latent space models, and it’s given by:
$\begin{split}\mathbf{z}_{i}&\in\mathbb{R}^{d}\\ \mathbf{z}_{i}&\overset{iid}{\sim}\text{MVN}_{2}(\mathbf{0},\gamma\mathbf{I}_{2})\\ Y_{ij}|\mathbf{z}_{i},\mathbf{z}_{j}&\sim\text{Bernoulli}(p_{ij}),i\neq j\\ p_{ij}&=\tau\exp\left\{-\frac{\left[d_{\mathbb{R}^{d}}(\mathbf{z}_{i},\mathbf{z}_{j})\right]^{2}}{2\phi}\right\}\end{split}$ (14)
where $d$ is the dimensional of the underlying Euclidean space and in this study, we set $d=2$. $\gamma>0$ and is a parameter for the multivariate normal distribution, $\tau\in[0,1]$ and $\phi>0$, where $\tau$ controls the sparsity of the network and $\phi$ relates the probability of having an edge between two nodes with their distances in the latent space.
Rastelli et al. [2016] has also shown that both expected average degree and clustering coefficient have explicit forms for Gaussian LPM, and are given by:
$\bar{k}$ $=(N-1)\tau\left\{\frac{\phi}{2\gamma+\phi}\right\}^{\frac{d}{2}}$ (15)
$C$ $=\tau\left(\frac{\gamma+\phi}{3\gamma+\phi}\right)^{\frac{d}{2}}$ (16)
3.1.2 Hyperbolic Geometry
The underlying geometry for latent space models is not limited to Euclidean geometry. Krioukov et al. [2010] proposes a framework that assumes that the underlying geometry for complex networks is hyperbolic geometry. The hyperbolic latent space model is given by:
$\begin{split}\mathbf{z}_{i}&=(r_{i},\theta_{i})^{\prime}\in\mathbb{H}^{d}\\ r_{i}&\overset{iid}{\sim}\text{Uniform}(0,R)\\ \theta_{i}&\overset{iid}{\sim}\text{Uniform}(0,2\pi)\\ Y_{ij}|\mathbf{z}_{i},\mathbf{z}_{j}&\sim\text{Bernoulli}(p_{ij}),i\neq j\\ logit(p_{ij})&=R-d_{\mathbb{H}^{d}}(\mathbf{z}_{i},\mathbf{z}_{j})\end{split}$ (17)
where $d$ is the dimensional of the underlying hyperbolic space and in this study, we set $d=2$. $R$ is the intrinsic radius, which controls how much latent space is being used. The latent positions are only sampled within disks of intrinsic radius $R$.
3.2 Network measures
The following network summary measures are used in this study:
  • Expected Average Degree $\bar{k}$, the expected number of connected nodes that each node has.
  • Density $den(\mathbf{Y})$: the proportion of all possible edges observed in a network.
  • Clustering coefficient $C$: if $n_{i}$ and $n_{j}$ are connected, and if $n_{j}$ and $n_{k}$ are connected, the clustering coefficient is the probability that $n_{i}$ and $n_{k}$ are connected.
3.3 Simulation details
For hyperbolic networks, based on Krioukov et al. [2010], if the network size is $N$, set $R=2\log\left(\frac{8n}{\pi\bar{k}}\right)$, where $\bar{k}$ is the target average degree and in this simulation study, by varying $\bar{k}$, we are able to generate networks for different densities. For GLPM, we choose $\gamma=1$ and $\phi=2$, and varying $\tau$ to get networks of various densities.
To evaluate the efficiency of Method 1, we tested networks of size 15 to 75 with densities ranging from 0 to 0.9. To compare all three methods, we first tested networks of size 15 to 45 with densities ranging from 0 to 0.9. Then, we focus on large and sparse networks whose sizes ranges from 50 to 200 and densities up to 0.2.
We apply classical MDS as implemented in the stats package in R. For MDS in hyperbolic space, we rely on the hydra package in R Keller-Ressel and Nargang [2019].
3.4 Results
Efficiency of Method 1
As shown in Figure 2, we notice that as network sizes get larger, the hyperbolic networks tends to be correctly identified, and when network size is larger than 60, almost all hyperbolic networks are correctly categorized as hyperbolic. For smaller hyperbolic networks, the probability of choosing correct hyperbolic geometry gets larger as the density of the network gets larger. However, this method doesn’t work well for Euclidean networks, and larger and/or denser networks tends to be misclassified as hyperbolic network.
(a) Hyperbolic Networks
(b) Euclidean Networks
Figure 2: The Percentage of Networks that MDS Classified as Hyperbolic
Compare Three Methods
We adopt the language of diagnostic tests. Let sensitivity indicate how well a method correctly identifies hyperbolic networks, and specificity indicate how well a method correctly identifies Euclidean networks.
First, we focus on smaller networks with a wide range of densities.
N Density Method 1 Method 2 Method 3
Sensitivity Specificity Sensitivity Specificity Sensitivity Specificity
15 0-0.2 0.4516 0.7593 0.1936 0.9537 0.2273 1
0.2-0.4 0.6905 0.3908 0.184 0.937 0.163 0.9474
0.4-1 0.714 0.2273 0.0828 0.9221 0.1037 0.9
30 0-0.2 0.8889 0.2158 0.3241 0.8633 0.2222 0.8861
0.2-0.4 0.9681 0.0677 0.4867 0.7961 0.3032 0.8469
0.4-1 0.9457 0.0581 0.2403 0.8968 0.2327 0.8903
45 0-0.2 0.9823 0.0072 0.6372 0.8129 0.354 0.8208
0.2-0.4 1 0.0242 0.7247 0.6667 0.5899 0.7621
0.4-1 0.9962 0.013 0.3277 0.8571 0.3089 0.8896
Table 1: Comparison of Three Proposed Methods on Smaller Networks
In general, for smaller networks, Method 1 could identify hyperbolic networks very well but, at the same time, tend to misclassify the majority of Euclidean networks as network getting larger and/or denser. Method 2 and Method 3 are more conservative and less likely to falsely categorize Euclidean networks as hyperbolic networks but misclassify the hyperbolic networks about more half the time.
Then, we focused on a particular group of networks - large and sparse networks. In this case, only networks of densities between 0 and 0.2 will be considered.
N Method 1 Method 2 Method 3
Sensitivity Specificity Sensitivity Specificity Sensitivity Specificity
50 0.9933 0.0333 0.6667 0.77333 0.4200 0.7750
100 1 0 0.94 0.6133 0.6133 0.6871
150 1 0 0.98 0.58 0.7333 0.6933
200 1 0 0.9867 0.7 0.84 0.7933
Table 2: Comparison of Three Proposed Methods on Large and Sparse Networks
From Table 2, it is clear that as the networks become larger, when network size is larger than 100, Method 1 categorizes all simulated networks, hyperbolic and GLPM, as hyperbolic networks and loses the effect of distinguishing the underlying geometry. Both Method 2 and Method 3 show an improvement in distinguishing between two methods for large and sparse network, which is more common in the real-world datasets. Method 2 has better sensitivity, that is good at correctly identifying hyperbolic networks, while Method 3 has a better balance between sensitivity and specificity.
4 Real Data
4.1 Karate Club
Papamichalis et al. [2022] use Zachary’s Karate club network as an example, compare the quality of MDS estimates when the manifold is assumed to be hyperbolic and Euclidean, and conclude that a model with a hyperbolic latent space better suits the data. Using the same package hydra as mentioned in this paper, we got $S_{\mathbb{R}^{2}}(\mathbf{Y})=24.65$ and $S_{\mathbb{H}^{2}}(\mathbf{Y})=18.20$. The stress from hyperbolic space is smaller, indicating a higher quality of embedding and shows that hyperbolic geometry would be better to model the data. Using Method 2 and Method 3, we get p-values of 0.0964 and 0.1468 respectively, which lead to the conclusion that Euclidean is better, as shown in Figure 3.
(a) Method 2, p-value: 0.0964
(b) Method 3, p-value: 0.1468
Figure 3: The Percentage of Networks that MDS Classified as Hyperbolic
4.2 Other Datasets
We have also applied these method on other datasets, a comparison of network characteristics and results from three different methods is presented below. The datasets used are from Nepusz et al. [2008](UKfaculty), Newman [2006](adjnoun), Lusseau et al. [2003](dolphin) and Kadushin [1995](ffe_friends).
(a) UKfaculty, n=81, dens.=0.178
Method 1: -19.32, Hyperbolic
Method 2: p-value=0, Hyperbolic
Method 3: N/A
(b) adjnoun, n=112, dens.=0.068
Method 1: -62.61, Hyperbolic
Method 2: p-value=0, Hyperbolic
Method 3: p-value=0, Hyperbolic
(c) dolphins, n=62, dens.=0.084
Method 1: 11.51, Euclidean
Method 2: p-value=0.99, Euclidean
Method 3: p-value=0.99, Euclidean
(d) ffe_friend, n=28, dens.=0.175
Method 1: -3.60, Hyperbolic
Method 2: p-value=0.20, Euclidean
Method 3: p-value=0.38, Euclidean
Figure 4: The average and standard deviation of critical parameters: Region R4
5 Discussion
We studied the effectiveness of using observed stress difference from MDS in determining the underlying geometry for latent space models through simulation study. It has been shown that when the stress difference is used alone, the geometry can’t be correctly identified especially for larger networks.
Two new methods based on the stress difference from MDS were proposed by introducing uncertainty with permutation test and bootstrapping. By studying the same sets of simulation networks, we showed that both methods outperform the original one, and work better on large and sparse networks which are commonly seen in real world.
The advantage of permutation test is that it retains the density of the observed network and doesn’t rely on the latent space models, but the structure of the network is not considered when generating the permuted networks. On the other hand, bootstrap method takes the structure of observed networks into account: if two nodes are currently connected, they are closer in latent space, thus more likely to be connected in the bootstrap networks. In this study, we used the Gaussian LPM as the Euclidean latent space model. However, for some networks, $\phi$ and $\tau$ could not be found through ad-hoc methods, thus bootstrapping could not be implemented.(Another thing about bootstrap method is that sometimes, it’s hard to get fully connected bootstrap network, and p-value can’t be found. haven’t test yet, but what if just get rid of the isolate nodes in new method?). This bootstrap method could be extended to other Euclidean models provided that the distribution of distance between nodes given their geodesic distance could be found theoretically or through Monte Carlo method.
However, both permutation and bootstrapping methods are computationally expensive since both classical MDS and hyperbolic MDS need to be applied to each of the permuted or bootstrapping networks.